1. Compute for the interest of a loan of P250,000 dated January 1, 2016 and payable on November 13, 2016 assuming the rate is 12% using the following: (20 points)

a. Actual Time-Exact Interest Method

b. Approximate Time-Exact Interest Method

c. Actual Time-Ordinary Interest Method

d. Approximate Time-Ordinary Interest Method


1
Expert's answer
2021-02-23T12:40:27-0500

a) 250000(12%)x 317/365=26054.79

b)250000x 0.12x 313/365= 25726.02

c) 250000x0.12x317/360= 26416.67

d) 250000x 0.12x 313/360= 26083.33